OpenStudy (anonymous):

verify the identity: (secθ+tanθ)(secθ-tanθ)=1

OpenStudy (anonymous):

could you this one? @.Sam.

sam (.sam.):

Use back the conversions, \[\tan(x)=\frac{\sin(x)}{\cos(x)}~~~~and~~~~\cot(x)=\frac{\cos(x)}{\sin(x)}\]

OpenStudy (anonymous):

and then do i cancel them out? @.Sam.

sam (.sam.):

Yeah, if you can, just cancel.

OpenStudy (anonymous):

thank you

sam (.sam.):

Some notes ;)

OpenStudy (anonymous):

:)